OpenStudy (lemanual):

Solve x divided by four equals eight . 4 −4 32 −32

OpenStudy (luigi0210):

\(\Large \frac{x}{4} =8 \) Multiply by 4: \(\Large 4*\frac{x}{4} =8*4 \) Now simplify it (: \(\Large x= 8*4 \)

OpenStudy (anonymous):

x/;4 = 8 multiply both sides by 4 4(x/4) = 4(8) x = 32

OpenStudy (lemanual):

Ty both :)